# Theory of two-component superfluidity of microcavity polaritons

### v1

@@ IV Tunable sound velocity
and,
$\beta({P})=\left[\omega_{\text{LP},{P}}^{2}-\omega_{\text{UP},{P}}^{2}\right]^{2}+\left(4G_{\text{LP-UP}}\right)^{2}\varepsilon_{\text{LP}}({P})\varepsilon_{\text{UP}}({P}).$ (34)
Substituting these into $\omega_{\text{LP},{P}}^{2}$ and $\omega_{\text{UP},{P}}^{2}$, we obtain
$\omega_{\text{LP},{P}}^{2}\approx\varepsilon_{\text{LP},0}^{2}+2G_{\text{LP}}\varepsilon_{\text{LP},0}+\frac{P^{2}}{m_{\text{LP}}}\left(\varepsilon_{\text{LP},0}+G_{\text{LP}}\right),$ (35)
$\omega_{\text{UP},{P}}^{2}\approx\varepsilon_{\text{UP},0}^{2}+2G_{\text{UP}}\varepsilon_{\text{UP},0}+\frac{P^{2}}{m_{\text{UP}}}\left(\varepsilon_{\text{UP},0}+G_{\text{UP}}\right).$ (36)
The cross-term $\beta({P})$ contributes
$\beta({P})\approx\text{constant terms}+P^{2}\left(\frac{G_{\text{LP-UP}}^{2}}{m_{\text{LP}}m_{\text{UP}}}\right).$ (37)
Thus, the sound velocity becomes,
$c_{j}=\sqrt{\frac{G_{\text{LP}}}{2m_{\text{LP}}}+\frac{G_{\text{UP}}}{2m_{\text{UP}}}+(-1)^{j-1}\sqrt{\left(\frac{G_{\text{LP}}}{2m_{\text{LP}}}-\frac{G_{\text{UP}}}{2m_{\text{UP}}}\right)^{2}+\frac{G_{\text{LP-UP}}^{2}}{m_{\text{LP}}m_{\text{UP}}}}}.$ (38)
In similar fashion the $j=1$ mode survives while the j = 2 mode leads to $c_{2}=0$, leaving us with the sound velocity of
$c_{1}=c_{s}=\sqrt{\left(\frac{G_{\text{LP}}}{m_{\text{LP}}}+\frac{G_{\text{UP}}}{m_{\text{UP}}}\right)}=\sqrt{U_{q}n_{0}\left(\frac{|X_{P}|^{2}}{m_{\text{LP}}}+\frac{|C_{P}|^{2}}{m_{\text{UP}}}\right)}.$ (39)
We are interested in how the sound velocity varies with the detuning and Rabi splitting. We plot these results below in Fig. 4.
Figure 4: The sound velocity as a function of the detuning in GaAs for different Rabi splittings $\Omega_{R}$. The parameters we kept constant are - total polariton concentration of $n_{0}=10^{8}$ cm-2, and a bare photon energy of 1.6 eV. We employed the definition of the sound velocity $c_{s}$ from Eq. (39).
Fig. 4 illustrates the sound velocity $c_{s}$ as a function of the detuning $\Delta_{0}$ for various values of the Rabi splitting $\Omega_{R}$. The graph reveals that at small detunings, particularly near $\Delta_{0}=0$, the sound velocity is the same regardless of the Rabi splitting. The definition of the effective masses of the polaritons in Eq. (70)-(71) show that at zero detuning, the masses do not vary with the Rabi splitting $\Omega_{R}$. Thus to investigate how the we may vary the sound velocity with the Rabi splitting, we select a detuning in the off-resonance regime. The results are shown below in Fig. 5.
Figure 5: The sound velocity as a function of Rabi splitting in GaAs for chosen detunings. Notably, the plots are symmetric for detunings equidistant from resonance, such that the curves for $+10$ meV and $-10$ meV are identical. The parameters employed are a total polariton concentration of $n_{0}=10^{8}\,\mathrm{cm}^{-2}$ and a bare photon energy of 1.6 eV. The sound velocity $c_{s}$ is determined using the definition in Eq. (39).
We note that in Fig. 5 the sound velocity is increased with increasing Rabi splitting. We may also analyze the differences in the sound velocity between the system with two-components versus one-component Bose-Einstein condensates, using the theoretical framework developed in previous sections. This can enlighten us on the advantages of considering our two-component condensate. The sound velocity in the one-component condensate of lower polariton only is given in Ref. berman2008 and is presented by Eq. (40), i.e.,
$c_{s}=\sqrt{\frac{U_{\text{eff}}n_{0}(r,T)}{M_{\text{eff}}}},$ (40)
In this notation, $U_{\text{eff}}$ is the strength of the effective interaction, $n_{0}(r,T)$ is the total polariton concentration, and $M_{\text{eff}}$ is the effective mass of the lower polariton. The plot in Fig. 6 highlights the distinct behavior of the sound velocity, $c_{s}$, as a function of detuning, $\Delta_{0}$, for both condensates.
Figure 6: (COLOUR ONLINE) The sound velocity as a function of the detuning for our two-component condensate of upper and lower polaritons, and the one-component condensate of lower polaritons only, in GaAs. We use the definition of the sound velocity from Eq. (40) and use the same lower polariton effective mass and the same total concentration of polaritons $n_{0}=10^{8}$ cm-2 in both. We have also included a plot of the lower polariton total polariton concentration as a fraction of the total concentration in both cases. We also used a Rabi splitting of $\Omega_{R}$ = 8 meV.
In Fig. 6, we compare the sound velocity of a polariton Bose gas with a one-component condensate (lower polaritons only) to that of a polariton Bose gas with a two-component condensate, in which the total polariton concentration $n_{0}$ is shared between upper and lower polaritons. Across all detunings, the polariton Bose gas with a two-component condensate exhibits a higher sound velocity. This arises because the two-component condensate benefits from contributions of both polariton branches, whose effective masses vary oppositely with detuning. As shown in Fig. 20(a), while the lower polariton becomes heavier at positive detuning, the upper polariton becomes lighter, and vice versa at negative detuning. This anti-symmetric behavior results in the polariton Bose gas with a two-component condensate displaying a consistently higher sound velocity. The largest difference occurs near zero detuning, where both branches are maximally mixed and their contributions most balanced.
@@ V Superfluidity
The density $\rho_{s}(T)$ of the superfluid is defined as $\rho_{s}(T)=\rho-\rho_{n}(T)$, where $\rho=m_{\text{UP}}n_{0}|C_{P}|^{2}+m_{\text{LP}}n_{0}|X_{P}|^{2}$ is the total 2D concentration of the system. Also, $\rho_{n}(T)$ is the density of the normal component. We define the normal component $\rho_{n}(T)$ in the usual way berman2022nanomaterials . We assume that the polariton system moves with velocity u, which means that the superfluid component has velocity u. At finite temperatures $T$, dissipating quasiparticles will emerge in this system. Since their density is small at low temperatures, one may assume that the gas of quasiparticles is an ideal Bose gas. We obtain the density of the superfluid component in the system following the procedure described in Ref. phosphoreneOGR . To calculate the superfluid component density, we begin by defining the mass current J for a Bose gas for quasiparticles in the frame of reference where the superfluid component is at rest by
$\textbf{J}=\int\frac{d^{2}P}{(2\pi\hbar)^{2}}\textbf{P}\left[f[\varepsilon_{1}(\textbf{P})-\textbf{P}\textbf{u}]+f[\varepsilon_{2}(\textbf{P})-\textbf{P}\cdot\textbf{u}]\right]\ ,$ (41)
where $f[\varepsilon_{1}(\textbf{P})]=\{\text{exp}[\varepsilon_{1}(\textbf{P})/(k_{B}T)]-1\}^{-1}$ and $f[\varepsilon_{2}(\textbf{P})]=\{\text{exp}[\varepsilon_{2}(\textbf{P})/(k_{B}T)]-1\}^{-1}$ are the Bose-Einstein distribution functions for the quasiparticles with the angle dependent dispersions $\varepsilon_{1}(\textbf{P})$ and $\varepsilon_{2}(\textbf{P})$ respectively. Expanding the expression under the integral in terms of $\textbf{P}\textbf{u}/k_{B}T$ and restricting ourselves to the first-order term, we obtain
$\textbf{J}=-\frac{\textbf{u}}{2}\int\frac{d^{2}P}{(2\pi\hbar)^{2}}\textbf{P}^{2}\left(\frac{\partial f[\varepsilon_{1}(\textbf{P})]}{\partial\varepsilon_{1}(\textbf{P})}+\frac{\partial f[\varepsilon_{2}(\textbf{P})]}{\partial\varepsilon_{2}(\textbf{P})}\right)\ .$ (42)
Using Eqs.  (41) and(42), along with the definition $\textbf{J}=\rho_{n}\textbf{u}$, we define
$\rho_{n}(T)=-\frac{1}{2}\int\frac{d^{2}P}{(2\pi\hbar)^{2}}\textbf{P}^{2}\left(\frac{\partial f[\varepsilon_{1}(\textbf{P})]}{\partial\varepsilon_{1}(\textbf{P})}+\frac{\partial f[\varepsilon_{2}(\textbf{P})]}{\partial\varepsilon_{2}(\textbf{P})}\right)\ .$ (43)
@@
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho c_{s}^{4}}{3\zeta(3)sk_{B}^{3}}\right]^{1/3}\ ,$ (44)

### v2

@@ III Spectrum of collective excitations
with interaction parameters
$G_{\text{LP}}$ $=g_{\ell\ell}\,n_{\text{LP},0},\quad g_{\ell\ell}=U_{0}|X_{0}|^{4},$ (30)
$G_{\text{UP}}$ $=g_{uu}\,n_{\text{UP},0},\quad g_{uu}=U_{0}|C_{0}|^{4},$ (31)
$G_{\text{LP-UP}}$ $=g_{\ell u}\,\sqrt{n_{\text{LP},0}n_{\text{UP},0}},\quad g_{\ell u}=U_{0}|X_{0}|^{2}|C_{0}|^{2}.$ (32)
The two-mode Bogoliubov spectrum for coupled condensates has been studied extensively in atomic systems Pindzola , where hybridization between branches arises from interspecies interactions. Here $n_{\mathrm{LP},0}$ and $n_{\mathrm{UP},0}$ denote the polariton densities in the lower and upper polariton branches, $U_{0}$ is the contact exciton-exciton interaction strength from Eq. (12), and $X_{0}$, $C_{0}$ are the Hopfield coefficients evaluated at $P=0$. We plot the resulting collective excitation spectrum in Figure 2 for GaAs quantum wells in a GaAs microcavity. While the Bogoliubov treatment of multi-component condensates is a well-established formalism, its application to the coupled LP–UP polariton system introduces a qualitatively new degree of tunability: both the effective masses and the interaction constants depend explicitly on the Hopfield coefficients. This dependence enables continuous tuning of the condensate’s collective properties through the cavity detuning and Rabi coupling, linking the underlying exciton–photon composition to experimentally observable superfluid characteristics.
Figure 2: The spectrum of collective excitations in GaAs as a function of momentum. The parameters used in the calculation are $\Omega_{R}=17$ meV, a bare photon of energy 1.6 eV, and a detuning $\Delta_{0}=0$ meV. The collective excitation spectrum is defined in Eq. (27).
Eq. (27) describes the collective excitation spectrum in the system with coupled lower- and upper-polariton condensates within the mean-field framework. The interaction coefficients satisfy $G_{\mathrm{LP\!-\!UP}}^{2}=G_{\mathrm{LP}}G_{\mathrm{UP}}$, and the $2\times 2$ interaction matrix becomes degenerate, so that the interaction matrix effectively reduces to a single dominant term. Physically, this condition implies that the two branches predominantly share a common total-density response, while their relative-density channel becomes weakly constrained. Within this simplified equilibrium treatment, the coupled condensates thus support one clearly defined linear (sound-like) excitation associated with the broken global $\mathrm{U}(1)$ symmetry and a second, low-energy mode that is only weakly dispersive. As shown in Figure 2, the $j=1$ branch corresponds to the collective sound-like mode with velocity $c_{1}$ from Eq. (35), while the $j=2$ branch may be interpreted as a relative-phase–like fluctuation between the LP and UP components. We emphasize that identifying this second mode with a distinct symmetry-breaking channel would require a full time-dependent analysis of the coupled order parameters. Such a treatment goes beyond the present static mean-field framework, which only captures equilibrium density responses. In this limit, the second branch should therefore be regarded as a soft collective excitation originating from interbranch coupling, rather than as a rigorously established Goldstone or Leggett mode Wouters2007 ; Leggett2001 . In the long-wavelength limit $P\to 0$, the excitation spectrum becomes linear, $\Omega_{j,P}\approx c_{j}|P|$, with sound velocities
$c_{j}=\sqrt{\frac{G_{\text{LP}}}{2m_{\text{LP}}}+\frac{G_{\text{UP}}}{2m_{\text{UP}}}+(-1)^{j-1}\sqrt{\left(\frac{G_{\text{LP}}}{2m_{\text{LP}}}-\frac{G_{\text{UP}}}{2m_{\text{UP}}}\right)^{2}+\frac{G_{\text{LP-UP}}^{2}}{m_{\text{LP}}m_{\text{UP}}}}}\ ,$ (33)
where $m_{\text{LP}}$ and $m_{\text{UP}}$ are the effective masses of the lower and upper polaritons, respectively. Eq. (33) shows how the Hopfield coefficients enter the collective spectrum through the renormalized couplings $G_{\text{LP}},G_{\text{UP}},$ and $G_{\text{LP-UP}}$. This formulation preserves the full detuning and Rabi splitting dependence of the excitonic and photonic fractions.
The parameterization of the condensate populations requires some care. The Hopfield coefficients $|X|^{2}$ and $|C|^{2}$ characterize the exciton–photon composition of an individual polariton mode, but they do not determine how the total polariton density is distributed between the lower (LP) and upper (UP) polariton branches. In realistic microcavity systems, this population balance is set by kinetic and reservoir processes—such as pumping, relaxation, and decay—or can be directly inferred from experiment. To capture this generally, we introduce a phenomologically motivated dimensionless population-split parameter $\alpha\in[0,1]$, which specifies the fraction of the total polariton density $n_{0}$ residing in the LP and UP branches as follows,
$n_{\mathrm{LP},0}=\alpha n_{0},\qquad n_{\mathrm{UP},0}=(1-\alpha)n_{0}.$ (34)
Once $\alpha$ (or equivalently $n_{\mathrm{LP},0}$ and $n_{\mathrm{UP},0}$) is specified, the Bogoliubov analysis proceeds self-consistently to yield the excitation spectrum, sound velocity, and critical temperature.
Our Bogoliubov analysis remains valid for any externally supplied $(n_{\text{LP},0},n_{\text{UP},0})$ from experiment or kinetic modeling. The subsequent sections explore how $c_{s}$ and $T_{c}$ depend on $\alpha$, providing testable signatures of two-component condensation.
IV Tunable sound velocity
Evaluating the expression for the sound-like branch ($j=1$ branch), and expanding, gives the effective sound velocity
$\boxed{c_{s}=\sqrt{U_{0}\!\left(\frac{|X_{0}|^{4}n_{\mathrm{LP},0}}{m_{\mathrm{LP}}}+\frac{|C_{0}|^{4}n_{\mathrm{UP},0}}{m_{\mathrm{UP}}}\right)}}.$ (35)
This form makes explicit how the sound velocity depends on the Hopfield coefficients, the effective masses of the polariton branches, and the population imbalance between them. We begin by visualising how the sound velocity as a function of detuning for various values of the population split parameter. In this work, the single-branch limits ($\alpha$ = 1 for LP-only and $\alpha$= 0 for UP-only) are modeled as realistic polariton condensates confined to one branch, retaining the Hopfield-dependent effective masses and interaction constants.
Figure 3: Comparison of sound velocity $c_{s}$ in a two-component polariton condensate: (a) dependence on detuning $\Delta_{0}$ at fixed $\Omega_{R}=8$ meV, and (b) dependence on Rabi splitting $\Omega_{R}$ at fixed $\Delta_{0}=5$ meV. In both cases the polariton density is $n=1\times 10^{11}~\text{cm}^{-2}$.
At zero detuning ($\Delta_{0}=0$), the lower and upper polariton fractions are equal ($|X_{0}|^{2}=|C_{0}|^{2}=1/2$), so their contributions to Eq. (35) balance, and the sound velocity converges to a single value for all population splits $\alpha$. This is because at zero detuning, the lower polariton and upper polariton effective masses, $m_{\text{LP}}$ and $m_{\text{UP}}$, stated analytically in Eq. (69)-(70), are equal. Away from resonance ($\Delta_{0}\neq 0$), the Hopfield coefficients $|X_{0}|^{2}$ and $|C_{0}|^{2}$ become asymmetric, and the relative condensate populations shift the overall mean-field energy. As shown in Figure 3(a), this causes the sound velocity to peak at different detunings depending on $\alpha$. For instance, when the condensate is predominantly lower polaritonic ($\alpha=0.75$), the stronger excitonic character ($|X_{0}|^{4}$ term) enhances the effective interaction, leading to a higher $c_{s}$ that peaks near $\Delta_{0}\!\approx\!5~\mathrm{meV}$. Increasing the bare Rabi splitting leads to higher sound velocities, reflected in larger $c_{s}$ at fixed $\Omega_{R}$ for greater $\alpha$. The figure illustrates this trend for discrete coupling strengths and is not intended to imply a continuous variation of $\Omega_{R}$. In Figure 3(b), a slightly positive detuning was chosen, so condensates with a greater lower-polariton fraction exhibit higher sound velocities, reflecting their stronger excitonic contribution to the interaction energy. We would also like to probe the sound velocity as a function of concentration for various total fixed polariton densities $n_{0}$ and $\alpha$ values in Figure (4).
Figure 4: Sound velocity $c_{s}$ in a two-component polariton condensate as (a) a function of the total polariton density $n_{0}$ for different values of the condensate population–split parameter $\alpha$, (b) and as a function of the condensate population–split parameter $\alpha$ for various fixed total polariton densities. The Rabi splitting is fixed at $\Omega_{R}=8~\text{meV}$ and the detuning at $\Delta_{0}=5~\text{meV}$.
The dependence of sound velocity on polariton density is shown in Figure (4)(a), where $c_{s}$ is plotted against $n_{0}$ on logarithmic scales for several values of the population-split parameter $\alpha$. The linear behavior in this log-log representation confirms the $c_{s}\propto\sqrt{n_{0}}$ scaling predicted by Eq. (35). At fixed density, condensates with higher LP fractions ($\alpha\to 1$) exhibit larger sound velocities due to the $|X_{0}|^{4}$ weighting of the LP self-interaction term, which is maximal at positive detunings where the LP branch is predominantly excitonic. Figure 4(b) shows the complementary view of the sound velocity $c_{s}$ versus $\alpha$ at fixed total densities $n_{0}$. At the chosen positive detuning ($\Delta_{0}=5~\mathrm{meV}$), the sound velocity grows monotonically with $\alpha$, interpolating smoothly between the pure UP limit and the pure LP limit for $n_{0}=1\times 10^{11}~\mathrm{cm}^{-2}$. This monotonic $\alpha$-dependence is characteristic of positive detunings, where the lower polariton has a larger excitonic component. For negative detunings, the situation is reversed: the upper polariton becomes more excitonic, leading to a corresponding decrease in $c_{s}$ with increasing $\alpha$.
@@ V Superfluidity
The density $\rho_{s}(T)$ of the superfluid is defined as $\rho_{s}(T)=\rho-\rho_{n}(T)$, where $\rho=m_{\text{UP}}n_{\text{UP},0}+m_{\text{LP}}n_{\text{LP},0}$ is the total 2D concentration of the system. Also, $\rho_{n}(T)$ is the density of the normal component. We define the normal component $\rho_{n}(T)$ in the usual way berman2022nanomaterials . We assume that the polariton system moves with velocity u, which means that the superfluid component has velocity u. At finite temperatures $T$, dissipating quasiparticles will emerge in this system. Since their density is small at low temperatures, one may assume that the gas of quasiparticles is an ideal Bose gas. We obtain the density of the superfluid component in the system following the procedure described in Ref. phosphoreneOGR . To calculate the superfluid component density, we begin by defining the mass current J for a Bose gas for quasiparticles in the frame of reference where the superfluid component is at rest by
$J=\int\frac{d^{2}P}{(2\pi\hbar)^{2}}P\left[f[\Omega_{1}(P)-Pu]+f[\Omega_{2}(P)-P\cdot]\right]\ ,$ (36)
where $f[\Omega_{1}(P)]=\{\text{exp}[\Omega_{1}(P)/(k_{B}T)]-1\}^{-1}$ and $f[\Omega_{2}(P)]=\{\text{exp}[\Omega_{2}(P)/(k_{B}T)]-1\}^{-1}$ are the Bose-Einstein distribution functions for the quasiparticles with the angle dependent dispersions $\Omega_{1}(P)$ and $\Omega_{2}(P)$ respectively. Expanding the expression under the integral in terms of $Pu/k_{B}T$ and restricting ourselves to the first-order term, we obtain
$J=-\frac{u}{2}\int\frac{d^{2}P}{(2\pi\hbar)^{2}}P^{2}\left(\frac{\partial f[\Omega_{1}(P)]}{\partial\varepsilon_{1}(P)}+\frac{\partial f[\Omega_{2}(P)]}{\partial\Omega_{2}(P)}\right)\ .$ (37)
Using Eqs. (36) and (37), along with the definition $J=\rho_{n}u$, we define
$\rho_{n}(T)=-\frac{1}{2}\int\frac{d^{2}P}{(2\pi\hbar)^{2}}P^{2}\left(\frac{\partial f[\Omega_{1}(P)]}{\partial\Omega_{1}(P)}+\frac{\partial f[\Omega_{2}(P)]}{\partial\Omega_{2}(P)}\right)\ .$ (38)
@@
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho c_{s}^{4}}{3\zeta(3)k_{B}^{3}}\right]^{1/3}\ ,$ (39)